a first number plus twice a second number is 22. Twice the first number plus the second totals to 23
Answer by robertb(5830)

let f = the first number, and s = the 2nd number.
f + 2s = 22,
2f + s = 23.
From the first equation, f = 22-2s. Substitute into the 2nd equation:
2(22-2s) + s = 23,
44 - 4s +s = 23,
-3s = -21,
s = 7, the second number.
f = 22-2s = 22 - 14 = 8, the first number.